OpenStudy (anonymous):

What is the total amount for an investment of $100 invested at 3.5% for 8 years and compounded continuously?

jimthompson5910 (jim_thompson5910):

A=Pe^(rt) A=100*e^(0.035*8) A=100*e^(0.28) A=100*1.32312981233744 A=132.312981233744 A=132.31 So the total amount is about $132.31

OpenStudy (anonymous):

a = a0 * e^(0.035*8) = 100 * e^0.28) = 132.31 dollars